Find the cartesian equation of the circleX= 2sin theta - 3 and Y= 2cos theta + 4

Grade:11

1 Answers

Arun
6 years ago
Dear Prabhu
 
X + 3 = 2 sin theta.       (1)
Y-4 = 2 cos theta.          (2)
Square 1 and 2 and add them
(X +3)² + (Y-4)² = 4 (sin²theta + cos² theta)
(X+3)² + (Y - 4)² = 4
This is the answer.
 
Regards
Arun (askIITians forum expert)